OpenStudy (anonymous):

Can someone please give me a hand? /: What is the equation, in standard form, of a parabola that models the values in the table? x => -2, 0, 4 f(x) => 4, -6, 70 a) y=3x^2+4x-6 b) y=4x^2+3x-6 c) y=6x^2+3x-4 d) y=-4x^2-3x+6 Any help would be amazingly helpful!

OpenStudy (anonymous):

@satellite73 ? @Hero ?

OpenStudy (anonymous):

@phi ? Anybody?

hero (hero):

Plug in the values

OpenStudy (anonymous):

How?

OpenStudy (phi):

x => -2, 0, 4 f(x) => 4, -6, 70 this table tells that when x is -2 you should get 4 when you use the equation I would start with x=0 (because 0 makes the arithmetic easy) what is f(0) in each of the choices A,B,C, and D ? f(0) is short-hand for : replace x with 0 in the equation a) y=3x^2+4x-6 b) y=4x^2+3x-6 c) y=6x^2+3x-4 d) y=-4x^2-3x+6 I'll do the first one a) y = 3*0^2 + 4*0 - 6 what is 0^2 (btw, 0^2 is short for 0*0) ? 0*0 is 0 and 3*0*0 is still 0 for the 2nd term 4*0 is ??? yes, 0 we have y= 0+0-6 = -6 so (a) could be the equation. But check the others. what do you get?

OpenStudy (anonymous):

b) y=-6

OpenStudy (anonymous):

c) y=-4

OpenStudy (anonymous):

d) y=6

OpenStudy (phi):

yes, and d) gives y = +6 the table x => -2, 0, 4 f(x) => 4, -6, 70 says f(0) should be -6. so it looks like it's either (a) or (b) I would check x= -2 in (a) and in (b) what is f(-2) using a) y=3x^2+4x-6

OpenStudy (anonymous):

22?

OpenStudy (phi):

no, let's do it in small steps. First, replace x with -2 (negative numbers are more tricky) y=3x^2+4x-6 replace x with -2: y = 3*(-2)^2 + 4*(-2) -6 is this what you started with ?

OpenStudy (anonymous):

Yes

OpenStudy (phi):

OK, what do you get for 3*(-2)^2 ? remember this means 3* -2 * -2

OpenStudy (anonymous):

12?

OpenStudy (phi):

yes -2*-2 is 4 (two negatives make a positive when multiplying) and 3*4 is 12 so we have y = 12 + 4*(-2) -6 can you finish ?

OpenStudy (anonymous):

y=12-8-6 y=-2?

OpenStudy (phi):

yes, and notice that the table says x => -2, 0, 4 f(x) => 4, -6, 70 we want to get 4, but we got -2. so (a) is out. that leaves (b) as the answer. If this were a test and you were short of time, I would go with (b). But if you have time, check (b) b) y=4x^2+3x-6 what is f(-2) with this equation ?

OpenStudy (anonymous):

y=4(-2)^2+3(-2)-6 y=16-6-6 y=4 ???

OpenStudy (phi):

yes looks good, 16-6 is 10 and 10-6 is 4 so (b) works. If we were gung-ho we could check f(4) and see that it gives 70 I would do that if you need the practice.

OpenStudy (anonymous):

Ok! Thank you so much!!!

OpenStudy (phi):

yw